nformation pertaining to Yekstop Corp.'s sales revenue is presented below:

November December January
Cash sales $ 101,000 $ 130,000 $ 83,000
Credit sales 293,000 455,000 239,000
Total sales $ 394,000 $ 585,000 $ 322,000

Management estimates that 4% of credit sales are eventually uncollectible. Of the collectible credit sales, 65% are likely to be collected in the month of sale and the remainder in the month following the month of sale. The company desires to begin each month with an inventory equal to 70% of the sales projected for the month. All purchases of inventory are on open account; 30% will be paid in the month of purchase, and the remainder paid in the month following the month of purchase. Purchase costs are approximately 60% of the selling prices.

Budgeted cash payments in December for November inventory purchases by Yekstop Corp. are:

Multiple Choice

  • $94,986.

  • $165,480.

  • $118,200.

  • $221,634.

  • $73,210.
Answer #1
November
Cost of goods sold ( Total sales * 60% = 394000 * 60% ) 236400
(+) Desired Ending inventory ( December total sales * 60% * 70% = 585000 * 60% * 70% ) 245700
(-) Beginning inventory ( November total sales * 60% * 70% = 394000 * 60% * 70% ) 165480
Budgeted purchases 316620
Budgeted cash payments in December for November inventory purchases = November budgeted purchases * ( 1 - 30% ) = 316620 * ( 1 - 30% )    221634